# TabImpute: Accurate and Fast Zero-Shot Missing-Data Imputation with a Pre-Trained Transformer

### v1

@@ 3.2 Synthetic Training Data Generation
We generate tens of millions of datasets with missing values to train our model through a two-step process: first, generating underlying data, then introducing missingness patterns on top.
@@ 3.2.1 Data Generation with Linear Factor Models
We generate data using a simple linear factor model (LFMs) (Bai & Ng, 2002). LFMs are commonly used in matrix completion literature to prove error bounds for matrix completion algorithms (Koren et al., 2009; Candes & Recht, 2012). This family of models generates a data matrix $Y\in\mathbb{R}^{m\times n}$ by assuming the data lies on or near a low-dimensional subspace. The simplest case generates the data matrix $Y$ as the inner product of two lower-rank latent factor matrices, $U\in\mathbb{R}^{m\times k}$ and $V\in\mathbb{R}^{n\times k}$, where $k\ll n,m$ is the rank:
@@
When training, we experimented with several classes of data-generating processes, including matrices from nonlinear factor models and structural causal models (SCM) similar to the ones used in TabPFN. However, we found that the model trained on linear factor models performed the best. See Tab. 5 in Sec. A.3 for results demonstrating this.
@@ 3.2.2 Missingness Patterns for Training and Evaluation
MCAR missingness means the probability of an entry being missing, defined as its propensity, is constant across entries and independent from any other randomness. The missingness indicators $M_{ij}$ are drawn i.i.d. from a Bernoulli distribution $M_{ij}\sim\mathrm{Bern}(p)$ across all $(i,j)\in[m]\times[n]$ for some constant $p\in(0,1)$. This is the simplest form of missingness, but is unrealistic (Van Buuren, 2012).
@@
For MAR missingness, the probability of an entry being missing depends only on the observed values $X$. In other words, the randomness in MAR can be explained by conditioning on observed factors. Additionally, every entry has a positive probability of being observed (i.e., $p_{ij}>0$). We simulate MAR through column-wise MAR, denoted Col-MAR: we choose several columns as predictor columns and use those values to mask entries in other columns. This is similar to the MAR approach taken in Jarrett et al. (2022).
@@
For the most complex missingness class, MNAR, the probability of an entry being missing can depend on unobserved factors. Note that MNAR patterns are significantly more difficult to handle systematically, often requiring specialized methods for a specific kind of MNAR pattern (Van Buuren, 2012). Due to the flexibility of our entry-wise featurization, TabImpute can produce imputations for these highly complex scenarios, even when columns are completely missing, such as in panel-data missingness patterns. HyperImpute was tested on two MNAR patterns in the Appendix of Jarrett et al. (2022), one where values are further masked after an MAR pattern and another where values outside a certain range are masked. We build on this work by testing on $11$ MNAR patterns (see Sec. A.2 and Tab. 6 for details). We implement a range of MNAR patterns to simulate plausible real-world scenarios. For example, we utilize the expressiveness of neural networks to create random propensity functions MNAR patterns, use bandit algorithms to induce column-adaptive missing patterns, simulate panel data missingness where some features are removed later, censoring where sensor readings fall outside a detectable range, and survey data artifacts like respondent polarization and skip-logic.
3.3 Training on Multiple Missingness Patterns
@@
We train our model to predict unobserved values under several missingness patterns simultaneously. For the final TabImpute model, we trained only on MCAR, NN-MNAR, and Self-Masking-MNAR because we found our model generalized to the other methods well without including them explicitly in training. We use the prior-data fitted negative log likelihood (NLL) loss proposed in Müller et al. (2022). Like other PFNs with continuous numerical output, we use the Riemann distribution output also proposed in Müller et al. (2022). Since we can generate an unlimited amount of synthetic data, we do not reuse any synthetic data and only do one gradient pass per batch of datasets. This allows our model to learn the underlying data-generating process and missingness mechanisms without risk of memorization. We use a learning rate of $0.0001$, a batch size of $64$, and train on around $25$ million synthetic datasets.
We use an adaptive algorithm to determine what proportion of missingness patterns we include in each batch as we train: Every $s$ gradient steps, we recalculate the proportion of each missingness pattern by: (i) create new batches of each missingness pattern, (ii) calculate the model’s loss value for each missingness pattern’s batch, and (iii) recalculate proportions by running softmax (Heaton, 2018, Pgs. 180-184) over the loss values. This procedure adaptively down-weights any missingness type our model performs well on and up-weights any missingness type our model performs poorly on. For our pre-trained TabImpute model, we set $s=50$ steps.
@@
Remark: We had initially attempted to train the model sequentially on one missingness pattern at a time, but found that the network always experienced catastrophic forgetting (McCloskey & Cohen, 1989) irrespective of learning rate (i.e., it forgot how to handle the previous missingness patterns).
3.4 Adaptive Ensembling
Our final method, TabImpute+, includes two layers of ensembling: (i) permutation-based pre-processing techniques and (ii) adaptive weighted average of TabImpute and EWF-TabPFN. Note that TabPFN’s regression class includes several ensembling techniques as well. The first level of ensembling performs random permutations of the rows and columns before running these permuted matrices through TabImpute, reversing the permutations, and then averaging the predictions across the runs. We used $4$ independent permutations for our final method.
[…]
@@ 4 Results on OpenML Datasets: MissBench
To evaluate TabImpute against other methods, we introduce MissBench: a missing-data imputation benchmark using $42$ OpenML (Vanschoren et al., 2013) tabular datasets with $13$ synthetic missingness patterns. For every dataset and missingness pattern, we test each method’s ability to impute masked values. The $42$ OpenML datasets span domains such as medicine, engineering, and education. The missingness patterns include $1$ MCAR pattern, $1$ MAR pattern, and $11$ MNAR patterns. We provide details for the MNAR patterns in Sec. A.2.
We test a suite of imputation methods on MissBench: column-mean imputation (Hawthorne & Elliott, 2005), SoftImpute (Hastie et al., 2015), MissForest (Stekhoven & Bühlmann, 2011), iterative chained estimators (ICE/MICE) (van Buuren & Groothuis-Oudshoorn, 2011; Royston & White, 2011), GAIN (Yoon et al., 2018), MIWAE (Mattei & Frellsen, 2019), an optimal transport-based method (Muzellec et al., 2020), and TabPFN’s imputation method (see Sec. 1).
@@
This normalization maps the best-performing method to 0 and the worst to 1. Finally, we define our Imputation Accuracy as 1Normalized RMSE1-$\text${Normalized RMSE}, where higher values indicate better performance.
@@
Tab. 1 presents results for each missingness pattern as well as overall performance. TabImpute+ achieves the best overall performance and for nearly all individual patterns. For completeness, we list the performance of methods not shown in the table in Sec. A.3 with non-normalized RMSE examples in Tab. 8. TabImpute+ performs best under high missingness conditions (Fig. 4, Sec. A.1), which is expected since it leverages generative pre-training rather than relying solely on available dataset information like discriminative methods.
@@
Table 1: Imputation Accuracy ± Standard Deviation by Missingness Pattern.
We train on the patterns above the dashed line.
Pattern TabImpute+ HyperImpute OT MissForest
MCAR 0.821 ± 0.157 0.804 ± 0.205 0.856 ± 0.114 0.867 ± 0.148
NN-MNAR 0.880 ± 0.126 0.762 ± 0.240 0.815 ± 0.139 0.819 ± 0.188
Self-Masking-MNAR 0.707 ± 0.279 0.717 ± 0.257 0.593 ± 0.246 0.666 ± 0.272
[2pt/2pt] Col-MAR 0.860 ± 0.178 0.819 ± 0.263 0.756 ± 0.262 0.776 ± 0.251
Block-MNAR 0.908 ± 0.168 0.873 ± 0.178 0.857 ± 0.152 0.860 ± 0.153
Seq-MNAR 0.905 ± 0.094 0.862 ± 0.193 0.892 ± 0.094 0.829 ± 0.203
Panel-MNAR 0.791 ± 0.329 0.515 ± 0.368 0.540 ± 0.374 0.526 ± 0.371
Polarization-MNAR 0.885 ± 0.137 0.620 ± 0.263 0.804 ± 0.125 0.559 ± 0.231
Soft-Polarization-MNAR 0.864 ± 0.176 0.705 ± 0.223 0.755 ± 0.195 0.670 ± 0.269
Latent-Factor-MNAR 0.887 ± 0.121 0.778 ± 0.231 0.844 ± 0.167 0.836 ± 0.152
Cluster-MNAR 0.873 ± 0.135 0.841 ± 0.183 0.828 ± 0.137 0.833 ± 0.141
Two-Phase-MNAR 0.855 ± 0.186 0.871 ± 0.200 0.809 ± 0.183 0.878 ± 0.123
Censoring-MNAR 0.594 ± 0.302 0.797 ± 0.262 0.599 ± 0.239 0.682 ± 0.249
Overall 0.833 ± 0.213 0.766 ± 0.259 0.765 ± 0.227 0.754 ± 0.248
Building on TabPFN.
The TabPFN team’s imputation extension predicts missing entries column-by-column using their pre-trained model. However, as shown in Fig. 1, this method has poor accuracy and slow runtime even on H200 GPU’s. This motivated us to introduce a new entry-wise featurization that achieves improved performance using TabPFN’s model, which we denote as EWF-TabPFN. We initially attempted fine-tuning TabPFN with our synthetic data but consistently encountered catastrophic forgetting, with performance degrading to match TabImpute, the model trained from scratch. Without access to TabPFN’s original training data, proper data augmentation proved immensely challenging. Instead, we elected to ensemble TabImpute and EWF-TabPFN as TabImpute+, leveraging their complementary performance across benchmark patterns (Tab. 2). Crucially, the performance scores in this table are re-normalized using only the results of these three methods. This means the minimum and maximum RMSE values used for normalization are determined from this smaller set, leading to different numerical scores than in Tab. 1, which was normalized across all 12 competing methods. This approach offers a clearer view of the relative performance and complementarity within our proposed family of models, highlighting which components excel at different missingness patterns. (See Sec. A.3 for other test results.)
Table 2: Imputation Accuracy ± Standard Deviation for Zero-Shot Methods.
Pattern TabImpute+ EWF-TabPFN TabImpute
MCAR 0.349 ± 0.468 0.324 ± 0.465 0.667 ± 0.477
NN-MNAR 0.551 ± 0.480 0.458 ± 0.488 0.487 ± 0.488
Self-Masking-MNAR 0.945 ± 0.214 0.952 ± 0.211 0.048 ± 0.216
[2pt/2pt] Col-MAR 0.799 ± 0.391 0.773 ± 0.414 0.220 ± 0.419
Block-MNAR 0.498 ± 0.474 0.437 ± 0.494 0.559 ± 0.501
Seq-MNAR 0.394 ± 0.482 0.363 ± 0.447 0.619 ± 0.492
Panel-MNAR 0.565 ± 0.476 0.568 ± 0.476 0.415 ± 0.499
Polarization-MNAR 0.775 ± 0.407 0.788 ± 0.411 0.214 ± 0.415
Soft-Polarization-MNAR 0.698 ± 0.422 0.791 ± 0.383 0.258 ± 0.439
Latent-Factor-MNAR 0.505 ± 0.490 0.467 ± 0.480 0.524 ± 0.505
Cluster-MNAR 0.428 ± 0.474 0.420 ± 0.487 0.571 ± 0.501
Two-Phase-MNAR 0.591 ± 0.475 0.560 ± 0.494 0.429 ± 0.501
Censoring-MNAR 0.891 ± 0.304 0.902 ± 0.282 0.098 ± 0.300
Overall 0.614 ± 0.468 0.600 ± 0.476 0.393 ± 0.487

### v3

@@ 3.2 Synthetic Training Data Generation
We generate around $1.9$ million matrices with missing values to train our model through a two-step process: first, generating underlying data, then introducing missingness patterns on top.
@@ 3.2.1 Data Generation with Linear Factor Models
We generate data using a simple linear factor model (LFMs) (linearfactor). LFMs are commonly used in matrix completion literature to prove error bounds for matrix completion algorithms (netflixprize; rechtexact). This family of models generates a data matrix $Y\in\mathbb{R}^{m\times n}$ by assuming the data lies on or near a low-dimensional subspace. The simplest case generates the data matrix $Y$ as the inner product of two lower-rank latent factor matrices, $U\in\mathbb{R}^{m\times k}$ and $V\in\mathbb{R}^{n\times k}$, where $k\ll n,m$ is the rank:
@@
When training, we experimented with several classes of data-generating processes (DGP), including matrices from nonlinear factor models and structural causal models (SCM) similar to the ones used in TabPFN. We found that training on SCM’s proved too computationally expensive, but were able to train a model on nonlinear factor models. We found that the model trained on linear factor models performed the best, as shown in Tab. 9 in Sec. A.3. We leave it as future work to explore other DGPs which might further enhance the underlying model.
@@ 3.2.2 Missingness Patterns for Training and Evaluation
MCAR missingness means the probability of an entry being missing, defined as its propensity, is constant across entries and independent from any other randomness. The missingness indicators $M_{ij}$ are drawn i.i.d. from a Bernoulli distribution $M_{ij}\sim\mathrm{Bern}(p)$ across all $(i,j)\in[m]\times[n]$ for some constant $p\in(0,1)$. This is the simplest form of missingness, but is unrealistic (van2012flexible).
@@
For MAR missingness, the probability of an entry being missing depends only on the observed values $X$. In other words, the randomness in MAR can be explained by conditioning on observed factors. Additionally, every entry has a positive probability of being observed (i.e., $p_{ij}>0$). We simulate MAR through column-wise MAR, denoted Col-MAR: we choose several columns as predictor columns and use those values to mask entries in other columns. This is similar to the MAR approach taken in jarrett2022hyperimpute.
@@
For the most complex missingness class, MNAR, the probability of an entry being missing can depend on unobserved factors. Note that MNAR patterns are significantly more difficult to handle systematically, often requiring specialized methods for a specific kind of MNAR pattern (van2012flexible). Due to the flexibility of our entry-wise featurization, TabImpute can produce imputations for these highly complex scenarios, even when columns are completely missing, such as in panel-data missingness patterns. HyperImpute was tested on two MNAR patterns in the Appendix of jarrett2022hyperimpute, one where values are further masked after an MAR pattern and another where values outside a certain range are masked. We build on this work by testing on $11$ MNAR patterns (see Sec. A.2 and Tab. 10 for details). We implement a range of MNAR patterns to simulate plausible real-world scenarios. For example, we utilize the expressiveness of neural networks to create random propensity functions MNAR patterns, use bandit algorithms to induce column-adaptive missing patterns, simulate panel data missingness where some features are removed later, censoring where sensor readings fall outside a detectable range, and survey data artifacts like respondent polarization and skip-logic.
3.3 Training TabImpute
@@
We train our model to predict unobserved values under several missingness patterns simultaneously. For the final TabImpute model, we trained only on MCAR missingness because we found our model generalized to the other patterns well without including them explicitly in training. In fact, we found that including other missingness patterns degraded overall performance. We discuss these results further below. We use the prior-data fitted negative log likelihood (NLL) loss proposed in muller2022transformers. Like other PFNs with continuous numerical output, we use the Riemann distribution output also proposed in muller2022transformers. Since we can generate an unlimited amount of synthetic data, we do not reuse any synthetic data and only do one gradient pass per batch of datasets. This allows our model to learn the underlying data-generating process and missingness mechanisms without risk of memorization. We use a learning rate of $0.0001$, a batch size of $16$, and train on around $1.9$ million synthetic datasets.
@@
Remark: We had initially attempted to train the model sequentially on one missingness pattern at a time, but found that the network always experienced catastrophic forgetting (mccloskey1989catastrophic) irrespective of learning rate (i.e., it forgot how to handle the previous missingness patterns). After this, we attempted to mix several missingness types together: MCAR, MAR, and Self-Masking-MNAR. However, we found that the model’s performance degraded overall because it focused only Self-Masking-MNAR. We attempted several mitigation techniques including pattern weight schedules and GradNorm chen2018gradnorm, a gradient normalization technique for multi-task learning, but found them unsuccessful in improving performance.
@@ 4 Results on OpenML Datasets: MissBench
To evaluate TabImpute against other methods, we introduce MissBench: a missing-data imputation benchmark using $42$ OpenML (OpenML2013) tabular datasets with $13$ synthetic missingness patterns. For every dataset and missingness pattern, we test each method’s ability to impute masked values. The $42$ OpenML datasets span domains such as medicine, engineering, and education. The missingness patterns include $1$ MCAR pattern, $1$ MAR pattern, and $11$ MNAR patterns. We provide details for the MNAR patterns in Sec. A.2.
We test a suite of imputation methods on MissBench: column-mean imputation (hawthorne2005imputing), SoftImpute (hastie2015matrix), MissForest (missforest2011), iterative chained estimators (ICE/MICE) (ice2011; royston2011multiple), GAIN (gain2018), MIWAE (mattei2019miwae), an optimal transport-based method (muzellec2020missing), and TabPFN’s imputation method (see Sec. 1). We also test imputing categorical variables via one-hot encoding in Sec. A.1 and report $R^{2}$ values in Tab. 8.
@@
This normalization maps the best-performing method to 0 and the worst to 1. Finally, we define Imputation Accuracy as 1Normalized RMSE1-$\text${Normalized RMSE}, where higher values indicate better performance.
@@
Tab. 1 presents results for each missingness pattern as well as overall performance. TabImpute achieves the best overall performance and for nearly all individual patterns. For completeness, we list the performance of methods not shown in the table in Tab. 6 and non-normalized RMSE values in Tab. 12. TabImpute performs best under high missingness conditions (Fig. 4, Sec. A.1), which is expected since it leverages generative pre-training rather than relying solely on available dataset information like discriminative methods. We also include column-wise 1-Wasserstein distance comparisons in Tab. 7 in Sec. A.3. In terms of the 1-Wasserstein distance, TabImpute and EWF-TabPFN do not perform as well as prior methods; however, we leave this for future work to investigate further.
Table 1: Imputation Accuracy ± Standard Error by Missingness Pattern.
We train on the pattern above the dashed line. The number of samples used to calculate standard error are the number of datasets: $42$.
Pattern TabImpute EWF-TabPFN HyperImpute MissForest
MCAR 0.882 ± 0.021 0.817 ± 0.024 0.804 ± 0.031 0.867 ± 0.022
NN-MNAR 0.900 ± 0.019 0.876 ± 0.020 0.760 ± 0.036 0.818 ± 0.029
Self-Masking-MNAR 0.703 ± 0.038 0.682 ± 0.042 0.689 ± 0.039 0.638 ± 0.040
Col-MAR 0.877 ± 0.030 0.855 ± 0.027 0.814 ± 0.040 0.773 ± 0.039
Block-MNAR 0.890 ± 0.026 0.905 ± 0.026 0.873 ± 0.027 0.860 ± 0.023
Seq-MNAR 0.920 ± 0.014 0.905 ± 0.015 0.864 ± 0.030 0.829 ± 0.031
Panel-MNAR 0.915 ± 0.025 0.711 ± 0.048 0.865 ± 0.033 0.912 ± 0.020
Polarization-MNAR 0.804 ± 0.024 0.879 ± 0.021 0.622 ± 0.039 0.567 ± 0.033
Soft-Polarization-MNAR 0.759 ± 0.045 0.864 ± 0.024 0.711 ± 0.034 0.667 ± 0.041
Latent-Factor-MNAR 0.891 ± 0.021 0.883 ± 0.018 0.775 ± 0.036 0.834 ± 0.023
Cluster-MNAR 0.903 ± 0.014 0.871 ± 0.021 0.839 ± 0.028 0.830 ± 0.021
Two-Phase-MNAR 0.886 ± 0.026 0.851 ± 0.029 0.865 ± 0.031 0.873 ± 0.020
Censoring-MNAR 0.710 ± 0.032 0.593 ± 0.045 0.786 ± 0.041 0.671 ± 0.038
Overall 0.849 ± 0.008 0.823 ± 0.009 0.790 ± 0.010 0.780 ± 0.009
Multiple imputation (MI) Metrics.
In multiple imputation, a dataset is imputed multiple times independently. These datasets can then be used to provide better confidence intervals of downstream quantities van2012flexible. Given two imputation methods that have the same imputation accuracy, one would prefer a method with higher uncertainty to not provide false precision on imputed values van2012flexible; li2015multiple; jolicoeur2024generating. We assess performance for multiple imputation by repeating imputation 5 times per dataset and calculating the median absolute deviation (MAD), minimimum RMSE, and average RMSE, shown in Tab. 2. While most imputation methods require another full run in order to create another imputation, TabImpute and other PFN-based methods output a distribution. This distribution is then sampled 5 times to create 5 different imputations, following the methodology of ForestDiffusion jolicoeur2024generating. Thus, multiple imputation for TabImpute uses the same runtime as a single imputation.
TabImpute provides better diversity than methods of comparable imputation accuracy (MissForest and HyperImpute). The poor precision of MissForest and HyperImpute is shown through their worse minimum RMSE values. Note that in Tab. 2, TabImpute has a higher avergage RMSE value then MissForest and HyperImpute. This is because it has a higher diversity in its estimate (the samples have a larger spread). When comparing just the median estimate, though, TabImpute has a better accuracy, as shown in the first row in Tab. 1. Thus, TabImpute provides practitioners with the best of both worlds: If they require multiple imputation for uncertainty estimates, TabImpute provides diverse estimates, and if they require a single good imputation, TabImpute provides the most accurate estimate.
Table 2: Multiple Imputation Metrics: Dataset-level MAD and RMSE Metrics ± Standard Error
per Imputed Entry Across 5 Samples (MCAR, p=0.4).
Method MAD ($\uparrow$) Min RMSE ($\downarrow$) Avg RMSE ($\downarrow$)
MICE 0.359 $\pm$ 0.002 0.264 $\pm$ 0.003 0.859 $\pm$ 0.005
MIWAE 0.267 $\pm$ 0.001 0.385 $\pm$ 0.005 0.927 $\pm$ 0.005
GAIN 0.200 $\pm$ 0.001 0.361 $\pm$ 0.004 0.789 $\pm$ 0.005
[2pt/2pt] TabImpute 0.199 $\pm$ 0.002 0.337 $\pm$ 0.004 0.722 $\pm$ 0.006
[2pt/2pt] ForestDiffusion 0.181 $\pm$ 0.001 0.291 $\pm$ 0.004 0.739 $\pm$ 0.005
OT 0.049 $\pm$ 0.000 0.440 $\pm$ 0.005 0.547 $\pm$ 0.005
HyperImpute 0.037 $\pm$ 0.001 0.420 $\pm$ 0.004 0.551 $\pm$ 0.005
MissForest 0.036 $\pm$ 0.000 0.434 $\pm$ 0.005 0.532 $\pm$ 0.005
SoftImpute 0.001 $\pm$ 0.000 0.577 $\pm$ 0.005 0.649 $\pm$ 0.005
Col Mean 0.000 $\pm$ 0.000 0.813 $\pm$ 0.005 0.813 $\pm$ 0.005
ICE 0.000 $\pm$ 0.000 0.606 $\pm$ 0.005 0.606 $\pm$ 0.005
Specialized TabImpute Models.
After training TabImpute on the MCAR pattern, we attempted to improve it by mixing in Col-MAR and Self-Masking-MNAR in the pre-training process. However, we found that the model consistently collapsed to only learning Self-Masking-MNAR, degrading performance on other patterns. This new model, though, performed far better on Self-Masking-MNAR than any other method, demonstrating the power of this architecture and pre-training to learn any pattern. We show the relative performance of three models trained on MCAR, Col-MAR, and Self-Masking-MNAR, respectively, in Tab. 3.
The model trained on Self-Masking-MNAR performs much better than the other models on this pattern. However, it does not generalize well to other patterns because the masks generated by the Self-Masking-MNAR pattern do not cover the types of masks generated by other patterns. On the other hand, when generating MCAR patterns millions of times, a small portion of these masks will look like Self-Masking-MNAR masks. Interestingly, the MCAR-trained model performed slightly better on Col-MAR patterns than the model trained on Col-MAR exclusively. One explanation for this is that the masks generated under MAR pattern are covered well by MCAR masks and MCAR masks produce a better underlying model. Note that the imputation accuracy scores are different for Tab. 1 and Tab. 3 because the MAR and Self-Masking-MNAR models are not included in the normalization in Tab. 1. However, the order of the imputation accuracy remains consistent.
Remark: We implemented an ensembling approach which weighted specialized model outputs by how well they predicted the observed values within a matrix, but found this approach to always favor the MCAR-trained model. We leave it for future work to determine a better method for ensembling or routing the specialized model predictions to better take advantage of specialized models.
@@
Table 3: Imputation Accuracy ± Standard Error by Missingness Pattern for Specialized Models
[…]
MCAR 0.938 ± 0.028 0.832 ± 0.024 0.020 ± 0.020
Self-Masking-MNAR 0.463 ± 0.067 0.329 ± 0.065 0.624 ± 0.072
Col-MAR 0.899 ± 0.042 0.811 ± 0.050 0.081 ± 0.041
Overall 0.766 ± 0.034 0.656 ± 0.035 0.243 ± 0.037